The circumference of a circle is 26 pi cm. Find the diameter, the radius, and the length of an arc of 170 degrees.
Answer by asuar010(338)

C=2*pi*r so 26pi=2*pi*r so r = 13... then diameter is 26 and to fond the length of an arc we use this formula 170/360*pi*13 and the answer is 19.28